Harris Company manufactures and sells a single product. A partially completed schedule of the companys total costs and costs per unit over the relevant range of 55,000 to 95,000 units is given below Required 1. Complete the schedule of the companys total costs and costs per unit as given in the relevant tab below. 2. Assume that the company produces and sells 85,000 units during the year at a selling price of $8.87 per unit. Prepare a contribution format income statement for the year. Complete this question by entering your answers in the tabs below. Required 1 Required 2 Complete the schedule of the companys total costs and costs per unit as given in the relevant tab below. (Round the per unit variable cost and fixed cost to 2 decimal places.) Units Produced and Sold 55,000 75,000 95,000 Total costs: Variable costs 132,000 410,000 542,000 Fixed costs Total costs Cost per unit: Variable cost Fixed cost Total cost per unit Required 1 Required 2
Answer #1

Requirement 1

Units produced and sold

55000

75000

95000

Total costs:

Variable costs

$ 1,32,000.00

$ 1,80,000.00

$ 2,28,000.00

Fixed costs

$ 4,10,000.00

$ 4,10,000.00

$ 4,10,000.00

Total costs:

$ 5,42,000.00

$ 5,90,000.00

$ 6,38,000.00

Cost per unit

Variable costs

$                2.40

$                2.40

$                2.40

Fixed costs

$                7.45

$                5.47

$                4.32

Total Cost per unit

$                9.85

$                7.87

$                6.72

Working

Units produced and sold

55000

75000

95000

Total costs:

Variable costs

132000

=132000/55000*75000

=132000/55000*95000

Fixed costs

410000

410000

410000

Cost per unit

Variable costs

=132000/55000

=180000/75000

=228000/95000

Fixed costs

=410000/55000

=410000/75000

=410000/95000

Total Cost per unit

=2.4+7.45

=2.4+5.47

=2.4+4.32

Requirement 2

Contribution margin income statement

Sales revenue

$ 7,53,950.00

Less: Variable cost (2.40 x 85000)

$ 2,04,000.00

Contribution margin

$ 5,49,950.00

Fixed cost

$ 4,10,000.00

Operating profit

$ 1,39,950.00